Two angles are complementary. The second angle measures 32 degrees less than the first angle. What is the measure of the first angle?
Answer by ptaylor(2198)


Two angles are complementary if they add up to 90 degrees.
Let x=first angle
Then x-32=second angle
So:
x+(x-32)=90 get rid of parens
x+x-32=90 add 32 to both sides
2x=122 divide both sides by 2
x=61 degrees ---------------------------first angle
x-32=61-32=29 degrees----------------------second angle
CK
61+29=90
also 61-32=29
29=29

Hope this helps----ptaylor